# From Data Behavior to Code Analysis: A Multimodal Study on Security and Privacy Challenges in Blockchain-Based DApp

### v1

@@ 2.2. DApp
A DApp(Izaguirre Diaz,, 2024) represents the integration of traditional applications (APPs) with blockchain technology, typically operating on a peer-to-peer (P2P) network. It serves as an enhancement and extension of conventional applications, with the key distinction being its decentralized nature. In DApp, participant information is either anonymous or protected, and operations are conducted through nodes on a peer-to-peer network. Smart contracts provide the foundational framework for decentralization, enabling trustless interactions between participants. From a practical perspective, a DApp can be succinctly conceptualized as a combination of smart contracts and traditional applications(Li et al.,, 2025), where smart contracts establish the prerequisites for decentralization. Structurally, DApp involves interactions between a front-end interface and users, as well as between smart contracts and the blockchain. This makes DApp publicly accessible programs that operate transparently on a network, leveraging blockchain’s inherent properties of immutability and traceability.
@@ 3.4.2. Model
An LSTM unit consists of an input gate $i_{t}$, an output gate $o_{t}$, a forget gate $f_{t}$, and a cell state $C_{t}$, allowing the unit to remember values at any time and control information flow. As shown in Figure.5.
@@
In this context,$C_{t}$ represents a new candidate vector. The $Sigmoid$ function $\sigma$ and the hyperbolic tangent function $tanh$ are activation functions used within the LSTM unit. The symbol $\odot$ denotes matrix multiplication and element-wise multiplication. These functions share similar equations but differ in their parameter matrices $W$. Since standard LSTM cannot capture future information in a sequence, a bidirectional LSTM layer is added to address this limitation.
To highlight the importance of certain output results for vulnerability detection, we introduced an Attention Mechanism, resulting in the BLSTM-ATT sequential model. For instance, for important words in lines of code (e.g., call.value), we use the Attention Mechanism to assign weights, which can be formalized as:
@@
We input word2vec-generated feature vectors into the BLSTM-ATT sequential model to learn model parameters. This involves calculating gradients and updating parameters during backpropagation. Once training is complete, the trained model is used for reentrancy detection.
@@
Formally, we use a $softmax$ classifier to predict the label $y^{*}$ of the contract snippet S. The detector takes the hidden state $h_{*}$ as input:
@@ 3.4.3. Model Training Details
In parameter settings, we use 10-fold cross-validation(Fushiki,, 2011) to select and train the optimal parameter values for reentrancy detection. We learn the model by optimizing binary cross-entropy loss. All experiments adopt the optimal gradient descent algorithm $Adam$(Kingma and Ba,, 2014). Our model searches for the learning rate $l_{r}$ in [0.0001, 0.0005, 0.001, 0.002, 0.005]. To prevent overfitting, we adjust the dropout rate $d_{r}$ searched in [0.2, 0.4, 0.6, 0.8]. The final parameters are set to: $l_{r}=0.02$, $d_{r}=0.2$, batch size $\beta=64$, and vector dimension $vm=300$.
@@ 3.5.2. Identity Privacy Protection Methods
Non-interactive Zero-Knowledge Proof(NIZK): A zero-knowledge proof (ZKP) is a cryptographic method that allows one to prove a statement without revealing any additional information. NIZK differs from ZKP in that it requires no interaction between the prover and verifier, making it suitable for anonymous and distributed message verification in blockchain systems. A formal definition of a Non-Interactive Zero-Knowledge (NIZK) proof system is as follows: Let ${(P,V)}$ denote a pair of probabilistic polynomial-time algorithms acting as the prover and verifier, respectively. For a language ${\mathcal{L}\subseteq\mathsf{NP}}$ (with a security parameter $k$), the tuple ${(P,V)}$ is called an NIZK proof system for $\mathcal{L}$ if it satisfies the following properties:

### v2

@@ 2.2. DApp
A DApp (Izaguirre Diaz,, 2024) represents the integration of traditional applications (APPs) with blockchain technology, typically operating on a peer-to-peer (P2P) (bukar2023peer, ; tariq2025peer, ) network. It serves as an enhancement and extension of conventional applications, with the key distinction being its decentralized nature. In DApp, participant information is either anonymous or protected, and operations are conducted through nodes on a peer-to-peer network. Smart contracts provide the foundational framework for decentralization, enabling trustless interactions between participants. From a practical perspective, a DApp can be succinctly conceptualized as a combination of smart contracts and traditional applications (Li et al.,, 2025), where smart contracts establish the prerequisites for decentralization. Structurally, DApp involves interactions between a front-end interface and users, as well as between smart contracts and the blockchain. This makes DApp publicly accessible programs that operate transparently on a network, leveraging blockchain’s inherent properties of immutability and traceability.
@@ 3.4.2. Model
An LSTM (hosseinzadeh2025comprehensive, ; alizadegan2025comparative, ) unit consists of an input gate $i_{t}$, an output gate $o_{t}$, a forget gate $f_{t}$, and a cell state $C_{t}$, allowing the unit to remember values at any time and control information flow. As shown in Figure.5.
@@
[…]
To highlight the importance of certain output results for vulnerability detection, we introduced an Attention Mechanism, resulting in the BLSTM-ATT (zhifeng2022comparison, ; tereshchenko2023vulnerability, ) sequential model. For instance, for important words in lines of code (e.g., call.value), we use the Attention Mechanism to assign weights, which can be formalized as:
@@
We input word2vec-generated feature vectors (abubakar2022sentiment, ) into the BLSTM-ATT sequential model to learn model parameters. This involves calculating gradients and updating parameters during backpropagation. Once training is complete, the trained model is used for reentrancy detection.
@@
Formally, we use a softmaxsoftmax classifier to predict the label $y^{*}$ of the contract snippet S. The detector takes the hidden state $h_{*}$ as input:
@@ 3.4.3. Model Training Details
In parameter settings, we use 10-fold cross-validation(Fushiki,, 2011) to select and train the optimal parameter values for reentrancy detection. We learn the model by optimizing binary cross-entropy loss. All experiments adopt the optimal gradient descent algorithm AdamAdam(Kingma and Ba,, 2014). Our model searches for the learning rate $l_{r}$ in [0.0001, 0.0005, 0.001, 0.002, 0.005]. To prevent overfitting, we adjust the dropout rate $d_{r}$ searched in [0.2, 0.4, 0.6, 0.8]. The final parameters are set to: $l_{r}=0.02$, $d_{r}=0.2$, batch size $\beta=64$, and vector dimension $vm=300$.
@@ 3.5.2. Identity Privacy Protection Methods
Non-interactive Zero-Knowledge Proof(NIZK): A zero-knowledge proof (ZKP) (wang2022non, ) is a cryptographic method that allows one to prove a statement without revealing any additional information. NIZK differs from ZKP in that it requires no interaction between the prover and verifier, making it suitable for anonymous and distributed message verification in blockchain systems. A formal definition of a Non-Interactive Zero-Knowledge (NIZK) (grilo2025distributed, ) proof system is as follows: Let (P,V){(P,V)} denote a pair of probabilistic polynomial-time algorithms acting as the prover and verifier, respectively. For a language $\mathcal{L}\subseteq\mathsf{NP}$} (with a security parameter $k$), the tuple (P,V){(P,V)} is called an NIZK proof system for $\mathcal{L}$ if it satisfies the following properties: